# FATE: A Formal Benchmark Series for Frontier Algebra of Multiple Difficulty Levels

### v2

@@ 1 Introduction
The emergence of reasoning models (Guo et al., 2025; OpenAI, 2025; Anthropic, 2025; Gemini Team, Google, 2025) has boosted the prospect that AI will assist in frontier mathematical research and produce rigorous proofs. Although modern mathematics has been successfully built upon the foundation of natural language proofs, complex proofs generated by current models remain unreliable. A major obstacle to improvement lies in verification: the more advanced a natural language proof is, the more it relies on limited human experts for rigorous verification, resulting in a process that is slow, unscalable, and error-prone. This creates a critical bottleneck for iterative development. In contrast, formal verification through proof assistants like Lean (Moura & Ullrich, 2021a) provides an automated, scalable, and reliable alternative for proof checking (see Appendix˜A for an introduction).
@@ 2 Related Works
Formal benchmarks enable reliable proof assessment. The miniF2F benchmark (Zheng et al., 2022) focuses on mathematical competitions, with similar efforts including Liu et al. (2023) and Tsoukalas et al. (2024). University-level formalization includes ProofNet (Azerbayev et al., 2022), hybrid datasets (Ren et al., 2025; Yu et al., 2025), specialized combinatorics benchmarks (Xiong et al., 2025; Liu et al., 2025), and the abstract algebra benchmark FATE-M (Shen et al., 2025). Hu et al. (2025) provides undergraduate-accessible problems extracted from formal projects by supplying all necessary lemmas and definitions. These are primarily implemented in Lean, building on Mathlib mathlib Community (2020).
@@ 3 The FATE Benchmark Series: Design and Curation
FATE focuses on abstract and commutative algebra, emphasizing proofs based on abstract properties and the study of algebraic structures instead of equation-solving or numerical computation, thereby reflecting the character of modern mathematics (see Section˜B.2.1 for common problem types). This area is particularly suitable for testing research-level reasoning due to its abstract and self-contained nature, which enables deep problem-solving without requiring extensive mathematical input from other fields. Lean’s existing algebra libraries provide a solid foundation for research-level questions, enabling one to attack advanced problems in this area with minimal additional formalization. Unlike benchmarks based on contest problems or introductory undergraduate material, FATE is designed to establish a difficulty progression toward modern mathematical research, structured as follows:
@@
To demonstrate that the mathematical difficulty is progressive and reaching the PhD qualifying exam level, we first describe the curation process in Section˜3.1, then focus on the mathematical content, followed by a description of formalization properties in Section˜3.2. Representative examples from each benchmark are provided in Section˜B.1.
@@ 3.1 Benchmark Curation
The curation process for FATE-H and FATE-X is highly demanding, as it requires experts with deep mathematical knowledge, specialists in Lean formalization, and extremely rare individuals with dual expertise. Figure˜2 illustrates the curation process.
@@
    Collection: We gathered and classified 400 candidate problems from the sources described above into two difficulty-based categories: 200 for FATE-H and 200 for FATE-X. The initial curation was conducted through a four-day collaborative effort involving nearly twenty experts—primarily postdocs and PhDs in pure algebra from top institutions—under the supervision of leading algebra researchers. This process ensured the final selection is mathematically profound, sound, well-balanced across key topics, and challenging for both human and machine reasoning.
@@
    Formalization: A subset of 200 problems (100 from each difficulty level) was formalized in Lean by five experienced formalization specialists, all contributors to Mathlib, including some involved in the ongoing formalization of Fermat’s Last Theorem (Buzzard, 2024). This effort was organized into five dedicated workshops, each involving over five hours of collaborative work.
@@
    Review: The formalizations were checked and corrected by two reviewers possessing both strong mathematical backgrounds and Lean expertise, each dedicating over 10 hours per benchmark. The accuracy was further verified with assistance from external leading experts in the Lean community.
@@ 3.2 Benchmark Characteristics
Problem distributions by mathematical topics in FATE-H and FATE-X are provided in Section˜B.2.2, including Figure˜3 visualizing the data. The charts illustrate a shift in focus from abstract algebra in FATE-H to advanced commutative algebra topics in FATE-X.
@@
Although a strong consensus exists among mathematicians regarding problem difficulty, it lacks objective, quantitative measures. We support this from three aspects:
  • Case Studies We present three examples (one from each benchmark) in Section˜B.3.1 involving similar mathematical topics with increasing difficulty. Analysis shows that these problems require: (1) direct application of a theorem as a linear deduction (FATE-M), (2) synthetic analysis of several direct results as integrative reasoning (FATE-H), and (3) exploration and analysis of new mathematical objects after synthesis as recursive structural analysis (FATE-X). This forms a clear difficulty progression.
  • Performance-based Metrics The following model performance data presented in Section˜B.3.2 reflect the difficulty progression across the three benchmarks: (1) the intermediate natural language accuracy during formalization of DeepSeek-R1 on FATE-M, FATE-H, and FATE-X decreases from nearly 94.7% to 71.0% to 33.0%; (2) the average output token length of reasoning models when solving natural language mathematical problems shows a clear increasing trend from M to H to X for each model.
  • Expert Assessment We conducted a questionnaire survey with ten professors from top institutions specializing in algebra (see Section˜B.4 for details). The results indicate that when comparing FATE to classical textbooks and ProofNet, experts consistently rated FATE significantly higher in difficulty, coverage, and originality. Regarding the suitability of FATE problems for PhD qualifying exams, 7/10 experts indicated that the problems were either directly appropriate for algebra/commutative algebra qualifying exams or notably more challenging than standard PhD qualifying exam questions.
@@
Due to its scope and difficulty, FATE-X uses mathematical definitions not yet in Mathlib. This introduces a distinctive formalization feature: new definitions are given before the problem statement. Consequently, LLMs must derive the necessary lemmas to use these definitions effectively. Moreover, to solve these challenging problems successfully, models are expected to discover mathematical phenomena, abstract them into useful lemmas, and, when needed, spontaneously formulate new definitions. This capability is critical for research-level mathematical formalization and problem-solving.
@@ 4 Experiments and Results
In Section˜4.1 and Section˜4.2, we detailed the experimental setup and baseline results on FATE. Observation of the model outputs indicated a two-stage generation pattern, where natural language precedes formalization. Consequently, Section˜4.3 and Section˜4.4 analyze the impact of this two-stage process on the models’ final accuracy separately. Section˜4.5 presents a comparative study between a general reasoning model and a theorem prover. Finally, Section˜4.6 provides a comprehensive discussion synthesizing these findings.
@@ 4.1 Experiment Setup
Our experiment included general reasoning models such as o3 (OpenAI, 2025), Gemini-2.5-Pro (Gemini Team, Google, 2025), Claude-4-Sonnet-thinking (Anthropic, 2025), and DeepSeek-R1 (Guo et al., 2025). We also tested state-of-the-art theorem provers, including DeepSeek-Prover-V2-671B (CoT), Kimina-Prover-72B, and Goedel-Prover-72B without self-correction mode. For both categories, we employed whole-proof generation, using pass@64 metric (Chen et al., 2021)-defined as successful if at least one correct proof is found within 64 independent attempts—with a maximum token limit of 64k. For detailed setup and prompt, see Sections˜F.1.1 and G.
During formal verification, we implemented multi-process parallelization using the Lean REPL [^2] https://github.com/leanprover-community/repl (a read-eval-print-loop for Lean4) to enable efficient large-scale evaluations. Every proof is rigorously checked by the Lean kernel to confirm it contains no sorry or compilation errors. Furthermore, to ensure semantic correctness and prevent unintended modifications, we employ string-matching checks to verify the accurate transcription of theorems and definitions. Any model-generated lemmas are also fully compiled to confirm their validity. A detailed methodology is available in Section˜F.1.2.
@@ 4.2 Benchmark Performance
Given the models’ two-stage process (natural language reasoning followed by formalization), we investigate the impact of each stage on final proof correctness in the following two subsections. For a discussion of the interaction between these stages, see section˜C.1.
@@ 4.3 Natural Language Reasoning Analysis
In this section, we first present a manual evaluation of the mathematical reasoning in the main experiment’s output in Section˜4.3.1, which reveals a significant gap between the correctness of the natural language portion and the final formalization. Based on this, we argue that formalization ability is the primary factor in Section˜4.3.3. To understand the gap between the intermediate result evaluated by us and the model’s real natural language mathematical ability, we also conduct an ablation study in Section˜4.3.2.
@@ 4.3.1 Manual Evaluation of Natural Language Proofs
The two-stage reasoning phenomenon observed in Section˜4.2 prompted us to question: do the models’ final formalization failures stem from errors in their initial mathematical reasoning, or from the translation process between natural language and formal language? To answer this, we organized a manual evaluation to independently assess the models’ natural language reasoning abilities under the original task settings.
Our mathematics experts evaluated the pass@1 correctness of the natural language proofs generated by DeepSeek-R1 and DeepSeek-Prover-V2 on the FATE-H and FATE-X. Manual evaluation criteria and procedures are detailed in Section˜F.2. We attribute errors to four main categories: Gap, Hallucination, Reasoning Problem, and No Progression. See Appendix˜C for definition and examples. The overall results of the evaluation are presented in Table˜2. Furthermore, we conducted a pass@16 evaluation for some problems and found that the models do exhibit diverse behaviors on the same problem; see Appendix˜C for details.
@@ 4.3.3 Analysis of Main Findings
    Primary Bottleneck: The Translation to Formal Language. First, for all models tested, intermediate natural language reasoning accuracy (pass@1) far exceeds final formalization accuracy (pass@64). The ablation study further suggests that the models’ mathematical ability when tasked only with generating a natural language proof provided with an informal statement is likely even higher than what this evaluation captured. In conjunction with our finding (see Section˜4.4) that formalization attempts are highly aligned with their preceding natural language reasoning, these results together show that the bottleneck is not the mathematical ability itself. Rather, the critical challenge lies in the translation and implementation of a correct natural language proof into an absolutely precise formal language.
@@
    Difference in Mathematical Abilities Between Models. At the level of natural language mathematical reasoning, the general-purpose reasoning model (DeepSeek-R1) behave significantly better than the specialized prover model (DeepSeek-Prover-V2). We will conduct a deeper comparative analysis of this phenomenon in Section˜4.5.
@@ 4.4 Formalization Error Analysis
This subsection classifies and quantifies common LLM formalization errors. Our human Lean experts analyzed proof attempts from DeepSeek-Prover-V2 on DeepSeek-R1 on the FATE-H benchmark, selected for its significant gap between natural language and formal language accuracy. After natural language evaluation in Section˜4.3.1, experts manually counted distinct formalization errors in mathematically correct but formally incorrect attempts.
These errors fall into four categories:
@@
For more details of these four categories, see Appendix˜D.
@@
For results in Table˜3, the denominator is the number of whole proofs generated by LLMs, which intermediate natural proof is manually judged to be correct. Among the four errors, Mathlib hallucinations and Lean proficiency issues were the most common errors, frequently recurring within almost every single proof attempt. In contrast, misalignment were remarkably infrequent. As DeepSeek-Prover-V2’s output lacks headers, we could not include its data for header-related issues. However, among other general capability issues, DeepSeek-Prover-V2 performed notably worse. Additionally, we found that only DeepSeek-Prover-V2 exhibited repetitive proof steps and misalignment.
A case study in Section˜D.2 revealed that the types of formalization errors on FATE-X were consistent with those on FATE-H. Notably, for problems involving new definitions, models rarely generated auxiliary lemmas to aid in proving.
@@
As demonstrated by the experimental results in Section˜4, the general-purpose reasoning model significantly outperforms the specialized proof-assistant model in terms of natural language success rate. To investigate the root cause of this performance gap, we conducted a detailed qualitative analysis of the reasoning processes of DeepSeek-V3 (Liu et al., 2024), DeepSeek-Prover-V2, and DeepSeek-R1-a series of models all post-trained from the same base model. More detailed experimental setup, results and case studies are provided in Appendix˜E. Here, we summarize the main conclusions:
@@
The intermediate natural language accuracy results of three models in Table˜8 confirm these observations: the DeepSeek-Prover-V2, which offers detail but lacks effective reflection, achieves a final success rate nearly identical to that of the formally reasoning DeepSeek-V3. Meanwhile, DeepSeek-R1, which possesses this capability, significantly outperforms both. Nevertheless, we must note that even DeepSeek-R1’s reasoning ability has its limitations when faced with truly complex reasoning environments.

### v3

@@ 1 Introduction
The emergence of reasoning models (Guo et al., 2025; OpenAI, 2025; Anthropic, 2025; Gemini Team, Google, 2025) has boosted the prospect that AI will assist in frontier mathematical research and produce rigorous proofs. Although modern mathematics has been successfully built upon the foundation of natural language proofs, complex proofs generated by current models remain unreliable. A major obstacle to improvement lies in verification: the more advanced a natural language proof is, the more it relies on limited human experts for rigorous verification, resulting in a process that is slow, unscalable, and error-prone. This creates a critical bottleneck for iterative development. In contrast, formal verification through proof assistants like Lean (Moura & Ullrich, 2021) provides an automated, scalable, and reliable alternative for proof checking (see Appendix A for an introduction).
@@ 2 Related Works
Formal benchmarks enable reliable proof assessment. The miniF2F benchmark (Zheng et al., 2022) focuses on mathematical competitions, with similar efforts including Liu et al. (2023) and Tsoukalas et al. (2024). University-level formalization includes ProofNet (Azerbayev et al., 2022), hybrid datasets (Ren et al., 2025; Yu et al., 2025), specialized combinatorics benchmarks (Xiong et al., 2025; Liu et al., 2025), and the abstract algebra benchmark FATE-M (Shen et al., 2025). Hu et al. (2025) provides undergraduate-accessible problems extracted from formal projects by supplying all necessary lemmas and definitions. These are primarily implemented in Lean, building on Mathlib (mathlib Community, 2020).
@@ 3 The FATE Benchmark Series: Design and Curation
FATE focuses on abstract and commutative algebra, emphasizing proofs based on abstract properties and the study of algebraic structures instead of equation-solving or numerical computation, thereby reflecting the character of modern mathematics (see Section B.2.1 for common problem types). This area is particularly suitable for testing research-level reasoning due to its abstract and self-contained nature, which enables deep problem-solving without requiring extensive mathematical input from other fields. Lean’s existing algebra libraries provide a solid foundation for research-level questions, enabling one to attack advanced problems in this area with minimal additional formalization. Unlike benchmarks based on contest problems or introductory undergraduate material, FATE is designed to establish a difficulty progression toward modern mathematical research, structured as follows:
@@
To demonstrate that the mathematical difficulty is progressive and reaching the PhD qualifying exam level, we first describe the curation process in Section 3.1, then focus on the mathematical content, followed by a description of formalization properties in Section 3.2. Representative examples from each benchmark are provided in Section B.1.
@@ 3.1 Benchmark Curation
The curation process for FATE-H and FATE-X is highly demanding, as it requires experts with deep mathematical knowledge, specialists in Lean formalization, and extremely rare individuals with dual expertise. Figure 2 illustrates the curation process.
@@
    Collection: We collected and classified 400 candidate problems into two difficulty-based categories: FATE-H and FATE-X, each containing 200 problems. The initial curation was conducted through a four-day collaborative effort involving nearly twenty postdocs and PhDs in pure algebra from top institutions, under the supervision of leading researchers. This process ensured the final selection is mathematically profound, well-balanced across key topics, and challenging for both human and machine reasoning.
@@
    Formalization: A subset of 200 problems (100 from each category) was formalized in Lean by five experienced specialists, all contributors to Mathlib, including some involved in the ongoing formalization of Fermat’s Last Theorem (Buzzard, 2024). This effort was organized into five dedicated workshops, each involving over five hours of collaboration.
@@
    Review: The formalizations were checked and corrected by two reviewers with strong mathematical backgrounds and Lean expertise, each dedicating over 20 hours. Accuracy was further verified with assistance from external experts in the Lean community.
@@ 3.2 Benchmark Characteristics
Problem distributions by mathematical topics in FATE-H and FATE-X are provided in Section B.2.2 and Figure 3. The charts illustrate a shift in focus from abstract algebra in FATE-H to advanced commutative algebra topics in FATE-X.
@@
Although a consensus exists among mathematicians regarding problem difficulty, it lacks objective measures. We support this from three aspects:
  • Case Studies We present three examples (one from each benchmark) in Section B.3.1 involving similar mathematical topics with increasing difficulty. Analysis shows that these problems require: (1) direct application of a theorem as a linear deduction (FATE-M), (2) synthetic analysis of several direct results as integrative reasoning (FATE-H), and (3) exploration and analysis of new mathematical objects after synthesis as recursive structural analysis (FATE-X). This forms a clear difficulty progression.
  • Human Performance Metrics We conducted a human experiment to quantitatively test the benchmark difficulty: Each participant, either a Ph.D. student or a postdoctoral researcher in algebra, was randomly assigned 5 non-overlapping natural language problems from FATE-H and FATE-X to solve in 2.5 hours. The participants achieved an accuracy of 73% on FATE-H and 21% on FATE-X, reflecting the significant difficulty progression across the benchmarks. The performance of LLMs also reflects a similar difficulty gradient, as detailed in Section B.3.2.
  • Expert Assessment We conducted a questionnaire survey with ten professors from top institutions specializing in algebra (see Section B.4 for details). The results indicate that when comparing FATE to classical textbooks and ProofNet, experts consistently rated FATE significantly higher in difficulty, coverage, and originality. Regarding the suitability of FATE problems for PhD qualifying exams, 7/10 experts indicated that the problems were either directly appropriate for algebra/commutative algebra qualifying exams or notably more challenging than standard PhD qualifying exam questions.
@@
Due to its scope and difficulty, FATE-X uses mathematical definitions not yet present in Mathlib. In total, 38% of its problems require new definitions, which are formalized before the problem statement. This subset of problems averages 2.4 new definitions each. These include advanced concepts from commutative algebra, such as local complete intersections and Gorenstein rings. Consequently, LLMs must derive the necessary lemmas to use these definitions effectively. Moreover, to solve these challenging problems successfully, models are expected to discover mathematical phenomena, abstract them into useful lemmas, and, when needed, spontaneously formulate new definitions. This capability is critical for research-level mathematical formalization and problem-solving.
@@ 4 Experiments and Results
In Sections 4.1 and 4.2, we detailed the experimental setup and baseline results on FATE. Observation of the model outputs indicated a two-stage generation pattern, where natural language precedes formalization. Consequently, Sections 4.3 and 4.4 analyze the impact of this two-stage process on the models’ final accuracy separately. Section 4.5 presents a comparative study between a general reasoning model and a theorem prover. Finally, Section 4.6 provides a comprehensive discussion synthesizing these findings.
@@ 4.1 Experiment Setup
Our experiment included general reasoning models such as o3 (OpenAI, 2025), Gemini-2.5-Pro (Gemini Team, Google, 2025), Claude-4-Sonnet-thinking (Anthropic, 2025), DeepSeek-R1 (Guo et al., 2025) and Qwen3 (Yang et al., 2025). We also tested state-of-the-art theorem provers, including DeepSeek-Prover-V2-671B (CoT), Kimina-Prover-72B, and Goedel-Prover-72B without self-correction mode. For both categories, we employed whole-proof generation, using pass@64 metric (Chen et al., 2021)-defined as successful if at least one correct proof is found within 64 independent attempts—with a maximum token limit of 64k. For detailed setup and prompt, see Sections F.1.1 and G.
During formal verification, we implemented multi-process parallelization using the Lean REPL [^2] https://github.com/leanprover-community/repl (a read-eval-print-loop for Lean4) to enable efficient large-scale evaluations. Every proof is rigorously checked by the Lean kernel to confirm it contains no sorry or compilation errors. Furthermore, to ensure semantic correctness and prevent unintended modifications, we employ string-matching checks to verify the accurate transcription of theorems and definitions. Any model-generated lemmas are also fully compiled to confirm their validity. A detailed methodology is available in Section F.1.2.
@@ 4.2 Benchmark Performance
Given the models’ two-stage process (natural language reasoning followed by formalization), we investigate the impact of each stage on final proof correctness in the following two subsections. For a discussion of the interaction between these stages, see section C.1.
@@ 4.3 Natural Language Reasoning Analysis
In this section, we first present a manual evaluation of the mathematical reasoning in the main experiment’s output in Section 4.3.1, which reveals a significant gap between the correctness of the natural language portion and the final formalization. Based on this, we argue that formalization ability is the primary factor in Section 4.3.3. To understand the gap between the intermediate result evaluated by us and the model’s real natural language mathematical ability, we also conduct an ablation study in Section 4.3.2.
@@ 4.3.1 Manual Evaluation of Natural Language Proofs
The two-stage reasoning phenomenon observed in Section 4.2 prompted us to question: do the models’ final formalization failures stem from errors in their initial mathematical reasoning, or from the translation process between natural language and formal language? To answer this, we organized a manual evaluation to independently assess the models’ natural language reasoning abilities under the original task settings.
Our mathematics experts evaluated the pass@1 correctness of the natural language proofs generated by DeepSeek-R1, DeepSeek-Prover-V2, Kimina-Prover, and Goedel-Prover-V2 on FATE-H and FATE-X. Manual evaluation criteria and procedures are detailed in Section F.2. We attribute errors to four main categories: Gap, Hallucination, Reasoning Problem, and No Progression. See Appendix C for definition and examples. The overall results of the evaluation are presented in Table 2. Furthermore, we conducted a pass@16 evaluation for some problems and found that the models do exhibit diverse behaviors on the same problem; see Appendix C for details.
@@
DeepSeek-Prover-V2 39.0% 3.0% 9.0% 0.0%
Goedel-Prover-V2 48.0% 2.0% 8.0% 0.0%
Kimina-Prover 35.0% 2.0% 3.0% 0.0%
@@ 4.3.3 Analysis of Main Findings
    Primary Bottleneck: The Translation to Formal Language. First, for all models tested, intermediate natural language reasoning accuracy (pass@1) far exceeds final formalization accuracy (pass@64). The ablation study further suggests that the models’ mathematical ability when tasked only with generating a natural language proof provided with an informal statement is likely even higher than what this evaluation captured. In conjunction with our finding (see Section 4.4) that formalization attempts are highly aligned with their preceding natural language reasoning, these results together show that the bottleneck is not the mathematical ability itself. Rather, the critical challenge lies in the translation and implementation of a correct natural language proof into an absolutely precise formal language.
@@
    Difference in Mathematical Abilities Between Models. At the level of natural language mathematical reasoning, the general-purpose reasoning model behave significantly better than the specialized prover models in Table 2. We will conduct a deeper comparative analysis of this phenomenon in Section 4.5.
@@ 4.4 Formalization Error Analysis
This subsection classifies and quantifies common LLM formalization errors. Our human Lean experts analyzed proof attempts from DeepSeek-Prover-V2 on DeepSeek-R1 on the FATE-H benchmark, selected for its significant gap between natural language and formal language accuracy. After natural language evaluation in Section 4.3.1, experts manually counted distinct formalization errors in mathematically correct but formally incorrect attempts. These errors fall into four categories:
@@
For more details of these four categories, see Appendix D.
@@
For results in Table 3, the denominator is the number of whole proofs generated by LLMs, which intermediate natural proof is manually judged to be correct. Among the four errors, Mathlib hallucinations and Lean proficiency issues were the most common errors, frequently recurring within almost every single proof attempt. In contrast, misalignment were remarkably infrequent. This suggests that a retrieval-augmented generation (RAG) system, which retrieves relevant theorems from Mathlib and provides accurate type information, could help improve formalization performance. As DeepSeek-Prover-V2’s output lacks headers, we could not include its data for header-related issues. However, among other general capability issues, DeepSeek-Prover-V2 performed notably worse. Additionally, we found that only DeepSeek-Prover-V2 exhibited repetitive proof steps and misalignment.
A case study in Section D.2 revealed that the types of formalization errors on FATE-X were consistent with those on FATE-H. Notably, for problems involving new definitions, models rarely generated auxiliary lemmas to aid in proving.
@@
As demonstrated by the experimental results in Section 4, the general-purpose reasoning model significantly outperforms the specialized proof-assistant model in terms of natural language success rate. To investigate the root cause of this performance gap, we conducted a detailed qualitative analysis of the reasoning processes of DeepSeek-V3 (Liu et al., 2024), DeepSeek-Prover-V2, and DeepSeek-R1-a series of models all post-trained from the same base model. More detailed experimental setup, results and case studies are provided in Appendix E. Here, we summarize the main conclusions:
@@
The intermediate natural language accuracy results of three models in Table 8 confirm these observations: the DeepSeek-Prover-V2, which offers detail but lacks effective reflection, achieves a final success rate nearly identical to that of the formally reasoning DeepSeek-V3. Meanwhile, DeepSeek-R1, which possesses this capability, significantly outperforms both. Nevertheless, we must note that even DeepSeek-R1’s reasoning ability has its limitations when faced with truly complex reasoning environments.